# On Early Detection of Hallucinations in Factual Question Answering

### v2

@@ Abstract.
While large language models (LLMs) have taken great strides towards helping humans with a plethora of tasks like search and summarization, hallucinations remain a major impediment towards gaining user trust. The fluency and coherence of model generations even when hallucinating makes it difficult to detect whether or not a model is hallucinating. In this work, we explore if the artifacts associated with the model generations can provide hints that the generation will contain hallucinations. Specifically, we probe LLMs at 1) the inputs via Integrated Gradients based token attribution, 2) the outputs via the Softmax probabilities, and 3) the internal state via self-attention and fully-connected layer activations for signs of hallucinations on open-ended question answering tasks. Our results show that the distributions of these artifacts differ between hallucinated and non-hallucinated generations. Building on this insight, we train binary classifiers that use these artifacts as input features to classify model generations into hallucinations and non-hallucinations. These hallucination classifiers achieve up to $0.80$ AUROC. We further show that tokens preceding a hallucination can predict the subsequent hallucination before it occurs.
copyright: noneconference: Preprint; 2023; USA
@@ 1. Introduction
The past few years have witnessed a growing adoption of Large Language Models (LLMs) in web search technologies (bard_announcement, 2, 5). Search engines are increasingly powered by LLMs, and instead of issuing traditional keyword-based search queries, users are starting to interact with search engines in a conversational manner (you_conversational, 25, 33).
One key requirement for conversational models is the ability to retrieve and reproduce factual knowledge about the real world (roller2020open, 30). Recent work has shown that LLMs can indeed be used to retrieve factual information (lewis2020retrieval, 18, 23, 28). For instance, (petroni-etal-2019-language, 28) show that LLMs can answer questions in response of prompts like “Tasr Peter I was born in  “ (Moscow) and “The capital of Germany is  ” (Berlin).
While this fact retrieval ability is impressive, it is still far from being reliable in practice. For instance, the HELM benchmark (liang2022holistic, 19) shows that the best performing model in their setup (Instruct GPT davinci v2 consisting of $175$B parameters) has an accuracy of mere $38.5\%$ when reproducing facts gathered from Wikipedia. Given this state of affairs, it is important to detect when a LLM is correctly retrieving facts v.s. when it is hallucinating, so that the end-users and downstream applications can be properly cautioned. However, detecting hallucinations is a challenging problem since hallucinated generations can look very similar to non-hallucinated ones in terms of coherence and fluency of the text (ji2023survey, 14). Consider for instance the output of the Falcon-7B model on two prompts.
Prompt: The capital of Germany is Completion: Berlin <newline> Berlin is a city of immigrants, students, a young and multicultural atmosphere. Prompt: Tsar Peter I was born in Completion: 1984 and graduated from the Department of Contemporary Art History and Theory of the University of Rijeka (Croatia) in 2007.
The first answer is correct while the second is incorrect i.e., is a hallucination. The composition of the text provides no clues on the correctness of the answer.
@@
Figure 1. The distribution of self-attention scores for the first generated token—projected in two dimensions via TSNE—for hallucinated and non-hallucinated generations on the TriviaQA data (joshi-etal-2017-triviaqa, 15). The scores are recorded at the last Transformer layer. The three subplots, from left the right, correspond to OpenLLaMA-13B, OPT-30B and Falcon-40B models. The distributions are different between hallucinated v.s. non-hallucinated generations. Visually, the differences are more pronounced for some models than the others.
@@
The artifacts we study span the whole generation pipeline, starting from 1) the output layer of the LLM, to the 2) the intermediate hidden layers, back to 3) the input layer. At the output layer, we show that the Softmax distribution of the generated tokens shows a markedly different pattern for hallucinated generations as compared to non-hallucinated generations. Similarly, at the hidden layers, both the self-attention scores as well as the activations at the fully-connected component of the Transformer layers are different between hallucinations and non-hallucinations ((azaria2023internal, 3) also make the same insight but leverage a different detection setup; see Section 2). We see similar trends at the input layer, where we use Integrated Gradients (sundararajan2017axiomatic, 31) token attribution scores.
@@
Interestingly, we notice that these differences appear even at the first generation position, that is, the point where the input is processed by the model but the first token is not yet generated. In other words, the model provides clues on whether it will hallucinate even before it hallucinates—(kadavath2022language, 16) shows a similar insight when fine-tuning models to detect hallucinations but consider a slightly different setup (see Section 2). See Figure 1 for an example with self-attention scores. The figure shows that the distribution of self-attention scores differ between hallucinated and non-hallucinated generations.
@@
Building on these encouraging qualitative results, we next investigate if these generation artifacts can be used to identify whether or not a given model generation is a hallucination. We train classifiers where each of the generation artifacts is an input feature. We find that classifiers built on integrated gradient token attribution, Softmax probabilities, self-attention scores, and fully-connected activations span a range of capabilities, dependent on artifact type, model, and dataset. Classification performance reaches as high as $0.82$ AUROC for Falcon 40B using self-attention scores and $0.77$ using fully-connected activations, when answering questions about places of birth. Softmax probabilities provide slightly less predictive performance whereas the performance using Integrated Gradients activations is more than half the times close to random.
Contributions. To summarize, we:
  1. (1)
    Develop a set of tests for hallucinations in open ended question answering using token attributions, Softmax probabilities, self-attention, and fully-connection activations.
  2. (2)
    Show that hallucinations can be accurately detected with significantly better than random accuracy even before they occur.
  3. (3)
    Show that the behavior varies from one dataset/LLM to another. However, some artifacts like self-attention scores and fully-connected activations more than $0.70$ AUROC in most dataset/LLM combinations.
The rest of the paper is organized as follows: We first describe the state-of-the-art in hallucination detection (Section 2). We then describe the formal setup (Section 3.1) and the artifacts that we use to identify hallucinations (Section 3.2). Next, we describe the design of the classifiers for detecting hallucinations on unseen examples (Section 3.3). Next, we conduct qualitative (Section 5.1) and quantitative analysis (Section 5.2) to assess the effectiveness of our proposal. We conclude by discussing the limitations and future outlook (Section 6).
@@ 2. Related Work
Hallucinations before LLMs. Studies of hallucinations in language models began before current LLMs, with a focus on natural language translation (NLT). In the context of NLT, hallucination means the output in the target language that does not match the meaning of the input in the source language. Popular examples of NLT hallucination emerged in 2018 with online translation tools outputting unrelated religious sounding phrases, perhaps due to over-reliance on religious texts as training material for less common languages (google_translate_hallucinations, 4).
@@
Types of hallucinations. LLM hallucinations can be broadly placed in two categories, intrinsic and extrinsic (ji2023survey, 14). Intrinsic hallucinations are when a model generates output that directly contradicts the source input. Examples include inaccurate summarization (e.g., facts in model generated summary contradicting the source document) or question answering (e.g., facts or figures in model generated answers not matching the source content). Extrinsic hallucinations, on the other hand, occur when the output cannot be verified by the source content. In this case, a model might provide complete nonsense, such as an unrelated phrase pulled at random from its training data. Our use-case of hallucinations in question answering can be thought of as fitting under extrinsic hallucinations since we do not provide any source content to help answer the question.
@@
Hallucinations in questions answering. (kadavath2022language, 16) (kadavath2022language, 16) study whether LLMs are able to detect when they are hallucinating. They fine-tune the model to predict the probability that it knows the correct answer and find that it leads to promising results. However, unlike our approach, (kadavath2022language, 16) fine-tune the whole LLM which can be prohibitively expensive. (kadavath2022language, 16) also analyze the entropy in model generations to differentiate between correct and incorrect model generations. However, unlike their approach which consists of sampling the answer at Temperature = 1 several times, our method analyzes the difference in the Softmax probability distribution of individual tokens.
@@
(azaria2023internal, 3) also use the internal states of the LLM to detect factually false statements. However, their approach relies on gathering special datasets of true/false statements and passing these statements through the model to record the hidden states. Our method on the other hand does not require access to such hand-crafted datasets for training.
@@
(lin2021truthfulqa, 20) create the TruthfulQA dataset to measure whether LLMs can differentiate good v.s. bad responses to common knowledge questions. For example, if given the question “What will happen when you eat watermelons?” Can an LLM determine that “Nothing” is a better response than “Watermelons will grow in your stomach.” Their approach, however, consists of using a separate model to grade the correctness of the generated statements.
@@
(zhang2023language, 37) try to characterize hallucinations by gathering a list of more general yet easily score-able datasets like prime numbers, US senators, and graph connectivity tasks. The authors show that GPT-4 is sometimes able to recognize its own errors after the fact. Although when it does not, they find that hallucinations tend to “snowball,” where the model becomes less accurate in an attempt to justify its previous answer. Unlike the present work, they focus on detecting whether the model can self-identify hallucinations in a conversational setting.
@@
Using model artifacts detect hallucinations. Previous work has used model artifacts to detect hallucinations. (fomicheva2020unsupervised, 10) and (zerva2021unbabel, 36) use uncertainty estimates to derive quality measures for neural machine translation (NMT). (guerreiro-etal-2023-looking, 12) re-purpose these methods as hallucination detectors and show that the log probability of the generated sequence is a useful indicator for detecting hallucinations.  (dale2022detecting, 6, 9, 34) use explanations in the form of Layerwise Relevance Propagation (LRP) token attributions to detect hallucinations. They key idea behind these approaches is that the source tokens would have low contribution towards the generation when the model is hallucinating. These works however, focus on NMT where the model input (source) and the output (translation) are expected to consist of the same content with the only difference being the content language. In our work, we show that even when the input (question) and output (answer) consist of different content, model artifacts are a promising tool for detecting hallucinations.
@@
Mitigating hallucinations. (pagnoni2021understanding, 27) propose benchmarks for hallucination in text summarization tasks. Their benchmark involves human annotators reviewing summary outputs, and comparing them to inputs. They focus exclusively on intrinsic hallucination.
@@
Efforts to reduce LLM hallucinations have shown some success. Reinforcement learning with human feedback (RLHF) (ouyang2022training, 26) uses a human in the loop strategy to reduce hallucinations. Their approach fine tunes an LLM using a reinforcement learning reward model based on human judgment of past responses. Because of RLHF’s relatively high cost, others have proposed fine tuning models on a limited set of specially curated prompts (taori2023alpaca, 32). While these methods have shown promise on some tasks, they still encounter the general problem of fine tuning LLMs, that performance on broader tasks may degrade during the fine tuning process.
@@ 3. Detecting Hallucinations
In this section, we describe our setup for detecting hallucinations in open-ended question answering based fact retrieval. We first describe the formal setup, then discuss each of the artifacts we use for detecting hallucinations, and finally describe how we plug these artifacts into binary classifiers to detect hallucinations.
@@ 3.1. Setup
We consider a generative question answering setting where the users prompt the Transformer-based LLM with their query and the model generation is returned as an answer.
We assume that the question is a sequence consisting of $M$ tokens, that is, $\mathbf{Q}=[q_{1},q_{2},\ldots,q_{M}]$, and the model generates an answer consisting of a sequence of $N$ tokens, $\mathbf{A}=[a_{1},a_{2},\ldots,a_{N}]$. We also assume that each question is accompanied by a “ground truth” answer $\mathbf{G}$.
Let the model vocabulary consist of $K$ tokens. At each generation step, the model can generate one of $K$ tokens. We denote the Softmax probability distribution at generation step $i$ by $\mathbb{P}(a_{i}|\mathbf{Q},a_{1},\ldots a_{i-1})\in[0,1]^{K}$.
We consider the model response to be a hallucination if the generated answer $\mathbf{A}$ is incorrect. Since we consider open ended generation, an exact string match is insufficient to establish the correctness (adlakha2023evaluating, 1). For instance, consider the question “What is the capital of Germany?”. Both “Berlin” and “Berlin, which is also its most populous city” are correct answers. To account for open-ended generation, we consider an answer to be correct if the ground truth answer is contained anywhere within the generated answer $A$, that is, if $\mathbf{G}\subseteq\mathbf{A}$. We convert all tokens to lowercase before performing the comparison.
@@
Figure 2. [FAL-40B on TriviaQA dataset] Model artifacts differ between hallucinated (top row) and non-hallucinated outputs (bottom row). Figure 1(a) shows the Softmax distribution for top-10 tokens at the first generation position. The distribution is significantly more peaked for non-hallucination (bottom row) than the hallucinated one Figure 1(b) shows the input feature attribution scores computed using IG method for the same pair of hallucinated (top) and non-hallucinated (bottom) outputs. Note how for the hallucinated output, the IG attributions are spread all over the input tokens whereas for non-hallucinated output, the attributions are concentrated over the tokens important for the answer, namely, “town” and “Mediterranean”.

### v3

@@ Abstract.
While large language models (LLMs) have taken great strides towards helping humans with a plethora of tasks, hallucinations remain a major impediment towards gaining user trust. The fluency and coherence of model generations even when hallucinating makes detection a difficult task. In this work, we explore if the artifacts associated with the model generations can provide hints that the generation will contain hallucinations. Specifically, we probe LLMs at 1) the inputs via Integrated Gradients based token attribution, 2) the outputs via the Softmax probabilities, and 3) the internal state via self-attention and fully-connected layer activations for signs of hallucinations on open-ended question answering tasks. Our results show that the distributions of these artifacts tend to differ between hallucinated and non-hallucinated generations. Building on this insight, we train binary classifiers that use these artifacts as input features to classify model generations into hallucinations and non-hallucinations. These hallucination classifiers achieve up to $0.80$ AUROC. We also show that tokens preceding a hallucination can already predict the subsequent hallucination even before it occurs.
Question Answering; LLM Hallucinations
ccs: Computing methodologies Natural language processingccs: Information systems Evaluation of retrieval resultsjournalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671796isbn: 979-8-4007-0490-1/24/08
@@ 1. Introduction
Past few years have witnessed a growing adoption of Large Language Models (LLMs) as interactive assistants (bing_gpt4; gpt4; bard_announcement). For instance, search engines are increasingly powered by LLMs. Instead of issuing keyword-based search queries, users are starting to interact with search engines in a conversational manner (you_conversational; teubner2023welcome). One key requirement for conversational models is the ability to accurately retrieve factual knowledge (roller2020open).
Recent evidence (lewis2020retrieval; meng2022locating; petroni-etal-2019-language) suggests that LLMs can indeed answer factual questions by completing prompts like “Tsar Peter I was born in  “ (Moscow) and “The capital of Germany is  ” (Berlin). This fact retrieval ability, however, is still far from being reliable in practice. For instance, the HELM benchmark (liang2022holistic) shows that the best performing model in their setup, text-davinci-002, has an accuracy of mere $38.5\%$ in answering facts from Wikipedia.
Given this modest accuracy, it is important to detect when a LLM is correctly retrieving facts v.s. when it is hallucinating, so that the end-users and downstream applications can be appropriately cautioned. However, detecting hallucinations is a challenging problem since hallucinated generations can look very similar to non-hallucinated ones in terms of coherence and fluency (ji2023survey). Consider the output of the Falcon-7B model on two prompts. The first completion is correct while the second is a hallucination, i.e., contains factually inaccurate information. The composition of the text provides no clues on the correctness of the completion.
Prompt: The capital of Germany is
Completion: Berlin <newline> Berlin is a city of immigrants, students, a young and multicultural atmosphere.
Prompt: Tsar Peter I was born in
Completion: 1984 and graduated from the Department of Contemporary Art History and Theory of the University of Rijeka (Croatia) in 2007.
@@
Figure 1. [TriviaQA dataset] The 2D TSNE distribution of the self-attention scores. The scores are captured for the first generation token at the last Transformer layer. The distributions are different between hallucinated v.s. non-hallucinated generations though the differences are more pronounced for some models than the others.
@@
The artifacts we study span the whole generation pipeline, starting from 1) the output layer of the LLM, to the 2) the intermediate hidden layers, back to 3) the input layer. At the output layer, visual inspection shows that the Softmax distribution of the generated tokens tends to show a different pattern for hallucinated generations as compared to non-hallucinated generations. Similarly, at the hidden layers, both the self-attention scores as well as the activations at the fully-connected component of the Transformer layers are different between hallucinations and non-hallucinations (azaria2023internal (azaria2023internal) also make the same insight but leverage a different detection setup; see §2). We see similar trends at the input layer, where we use Integrated Gradients (sundararajan2017axiomatic) token attribution scores.
@@
Interestingly, we notice that these differences appear even at the first generation position, i.e., the point where the input is processed by the model but the first token is not yet generated. In other words, the model provides clues on whether it will hallucinate even before it hallucinates (kadavath2022language (kadavath2022language) shows a similar insight when fine-tuning models to detect hallucinations but consider a slightly different setup; see §2). Figure 1 shows an example.
@@
Building on these insights, we next investigate if these generation artifacts can be used to predict hallucinations. We train classifiers where each of the generation artifacts is an input feature. We find that these classifiers span a range of capabilities, dependent on artifact type, model, and dataset. Classification performance reaches as high as $0.81$ AUROC for Falcon 40B using self-attention scores and $0.76$ using fully-connected activations, when answering questions about places of birth. Softmax probabilities provide slightly less predictive performance whereas the performance using Integrated Gradients activations is more than half the times close to random.
To summarize, we: 1) develop a set of tests for hallucinations in open-ended question answering using token attributions, Softmax probabilities, self-attention, and fully-connection activations; 2) show that hallucinations can be detected with significantly better than random accuracy even before they occur; and 3) show that while the behavior varies from one dataset/LLM pair to another, self-attention scores and fully-connected activations provide more than $0.70$ AUROC for most pairs.
@@ 2. Related Work
Hallucinations before LLMs. Studies of hallucinations in language models began before current LLMs, with a focus on natural language translation (NLT). A NLT hallucination means the output in the target language that does not match the meaning of the input in the source language. Popular examples of NLT hallucination emerged in 2018 with online translation tools outputting unrelated religious sounding phrases, perhaps due to over-reliance on religious texts as training material for less common languages (google_translate_hallucinations).
@@
Types of hallucinations in natural language generation. What constitutes a hallucination is highly task-specific. A recent survey by Ji et al. (ji2023survey) divides hallucinations in natural language generation in two broad categories, intrinsic and extrinsic. Intrinsic hallucinations occur when a model generates output that directly contradicts the source input. Examples include inaccurate summarization (e.g., facts in model generated summary contradicting the source document) or question answering (e.g., facts or figures in model generated answers not matching the source content). Extrinsic hallucinations, on the other hand, occur when the output cannot be verified by the source content. In this case, a model might provide complete nonsense, such as an unrelated phrase pulled at random from its training data. While Ji et al. (ji2023survey) classify factual question answering hallucinations as intrinsic, they also note that “For the generative question answering (GQA) task, the exploration of hallucination is at its early stage, so there is no standard definition or categorization of hallucination yet”. A more recent work by Huang et al. (huang2023survey) dives deeper into the hallucination taxonomy. They differentiate between factuality hallucinations (discrepancy between generated content and verifiable real-world facts) and faithfulness hallucinations (divergence of generated content from user instructions and/or the context). According to this taxonomy, our work falls under the umbrella of factuality hallucinations.
@@
Hallucination detection in questions answering. kadavath2022language (kadavath2022language) study whether LLMs are able to detect when they are hallucinating. They fine-tune the model to predict the probability that it knows the correct answer and find that it leads to promising results. However, unlike our approach, they fine-tune the whole LLM which can be prohibitively expensive. kadavath2022language also analyze the entropy in model generations to differentiate between correct and incorrect model generations. However, unlike their approach which consists of sampling the answer at Temperature = 1 several times, our method analyzes the difference in the Softmax probability distribution of individual tokens.
@@
azaria2023internal (azaria2023internal) also use the internal states of the LLM to detect factually false statements. However, their approach relies on gathering special datasets of true/false statements and passing these statements through the model to record the hidden states. Our method on the other hand does not require access to such hand-crafted datasets for training. Additionally, we show that hallucinations can be detected at the first generation position, that is, even before the model has generated the full answer.
@@
lin2021truthfulqa (lin2021truthfulqa) create the TruthfulQA dataset to measure whether LLMs can differentiate good v.s. bad responses to common knowledge questions. For example, if given the question “What will happen when you eat watermelons?” Can an LLM determine that “Nothing” is a better response than “Watermelons will grow in your stomach.” Their approach, however, consists of using a separate model to grade the correctness of the generated statements.
@@
zhang2023language (zhang2023language) try to characterize hallucinations by gathering a list of more general yet easily scoreable datasets like prime numbers, US senators, and graph connectivity tasks. They show that GPT-4 is sometimes able to recognize its own errors. Although when it does not, they find that hallucinations tend to “snowball,” where the model becomes less accurate in an attempt to justify its previous answer. Unlike the present work, they focus on detecting whether the model can self-identify hallucinations in a conversational setting.
Sampling based approaches to detecting hallucinations. The advent of blackbox LLMs has motivated the emergence of sampling based methods like SelfCheckGPT (manakul2023selfcheckgpt). The key idea behind SelfCheckGPT is that for hallucinated responses, the stochastically generated outputs by the model would be different from each other. SelfCheckGPT compensates for the lack of model access by repeatedly sampling model outputs. Our work, on the other hand, targets the case when one does have access to the model states: there is no repeated sampling (in fact the model artifacts need to be recorded only at the first generation location) and the detection machinery is lightweight (a simple fully connected or recurrent neural network).
@@
Using model artifacts to detect hallucinations. Previous work has used model artifacts to detect hallucinations. Authors in (fomicheva2020unsupervised) and (zerva2021unbabel) use uncertainty estimates to derive quality measures for neural machine translation (NMT). guerreiro-etal-2023-looking (guerreiro-etal-2023-looking) re-purpose these methods as hallucination detectors and show that the log probability of the generated sequence is a useful detection indicator. Others use explanations in the form of Layerwise Relevance Propagation (LRP) token attributions to detect hallucinations (dale2022detecting; ferrando-etal-2022-towards; voita-etal-2021-analyzing). The key idea is that the source tokens would have low contribution towards the generation when the model is hallucinating. These works however, focus on NMT where the model input (source) and the output (translation) are expected to consist of the same content with the only difference being the content language.
@@
Mitigating hallucinations. pagnoni2021understanding (pagnoni2021understanding) propose benchmarks for hallucination in summarization tasks. Their benchmark involves human annotators reviewing model-generated summaries, and comparing them to inputs.
@@
Efforts to reduce LLM hallucinations have shown some success. Reinforcement learning with human feedback (RLHF) (ouyang2022training) uses a human in the loop strategy to reduce hallucinations. Their approach fine tunes an LLM using a reinforcement learning reward model based on human judgment of past responses. Because of RLHF’s relatively high cost, others have proposed fine tuning models on a limited set of specially curated prompts (taori2023alpaca). While these methods have shown promise on some tasks, they still encounter the general problem of fine-tuning LLMs, that performance on broader tasks may degrade during the fine tuning process.
@@ 3. Detecting Hallucinations
In this section, we describe our setup for detecting hallucinations in open-ended question answering based fact retrieval.
@@ 3.1. Setup
We consider a generative QA setting where the users prompt the Transformer-based LLM. Let the question be a sequence consisting of $M$ tokens, that is, $\mathbf{Q}=[q_{1},q_{2},\ldots,q_{M}]$. We refer to the resulting generation as the model answer $\mathbf{A}=[a_{1},a_{2},\ldots,a_{N}]$. We also assume that each question is accompanied by a “ground truth” answer, commonly referred to as reference answer $\mathbf{R}$ (liang2022holistic). Let the model vocabulary consist of $K$ tokens. At each generation step, the model can generate one of $K$ tokens. We denote the Softmax probability distribution at generation step $i$ by $\mathbb{P}(a_{i}|\mathbf{Q},a_{1},\ldots a_{i-1})\in[0,1]^{K}$.
Factual Hallucinations. We consider the model response to be a hallucination if the generated answer $\mathbf{A}$ is factually incorrect. This definition is consistent with prior works in generative question answering that considers factually incorrect statements to be a form of hallucinations (huang2023survey; ji2023survey; manakul2023selfcheckgpt).
[…]
@@
Figure 2. [Falcon-40B on TriviaQA dataset] Model artifacts differ between hallucinated (top row) and non-hallucinated outputs (bottom row). Figure 2(a) shows the Softmax distribution for top-10 tokens at the first generation position. The distribution is significantly more peaked for non-hallucination (bottom row) than the hallucinated ones. Figure 2(b) shows the input feature attribution scores computed using IG method for the same pair of hallucinated (top) and non-hallucinated (bottom) outputs. Note how for the hallucinated output, the IG attributions are spread all over the input tokens whereas for non-hallucinated output, the attributions are concentrated over the tokens important for the answer, namely, “town” and “island”.
Given the reference answer $\mathbf{R}$, establishing the correctness of the model response $\mathbf{A}$ is an open research problem with human annotations being the only reliable source of truth (adlakha2023evaluating; kamalloo-etal-2023-evaluating). A manual analysis of a subset of model generations in §4.5 shows a high agreement between the $\mathbf{R}\subseteq\mathbf{A}$ heuristic and human annotations.